# A Survey of Cross-domain Graph Learning: Progress and Future Directions

### v1

@@ II-D Cross-domain graph learning
Definitions. Derived from cross-domain learning, CDGL aims to capture cross-domain knowledge from given source-domain graph data $\mathcal{G}=\{\mathcal{G}_{S_{1}},\dots,\mathcal{G}_{S_{i}}|i=1,\dots,m\}$ and then transfer the captured knowledge to learn a set of task functions (e.g., node classification, graph classification, etc.) on target-domain graph data $\mathcal{G}=\{\mathcal{G}_{T_{1}},\dots,\mathcal{G}_{T_{i}}|i=1,\dots,n\}$.
@@ II-D1 The Refinement of Cross-domain Scale
    Limited cross-domain graph learning refers to scenarios where graph samples from source and target domains share a set of similar or consistent attributes (i.e., from the same field). For example, NaP [35] divides the Facebook100 dataset—which contains social networks from 100 U.S. universities—into different domains, treating the social network of each university as a unique domain. The data from each domain share similar attributes, except for the university types. CDGL is then conducted across different universities. Similarly, SDGA [44] treats ACMv9, Citationv1, and DBLPv7—datasets constructed from ArnetMiner [87]—as distinct domains for cross-domain learning.
@@
    Conditional cross-domain graph learning refers to situations where the boundaries between source domains and graph domains are clearly defined, but their attributes can be transformed into a common format (e.g., text) and covered by a specific high-level field. For example, OFA [38] can convert cross-domain graph data into a text-attribute graph (TAG) format. However, it faces difficulties when applied to non-TAG domains, such as brain networks [60].
@@
    Open cross-domain graph learning refers to scenarios where graph data from both source and target domains can be arbitrary and do not need to satisfy specific conditions. For example, GCC [26] focuses on leveraging only the structural information from datasets across domains, using it to pre-train models and learn common structural patterns in graph data. GCOPE [21] presents a general framework for cross-domain graph pre-training that unifies the feature dimensions of graph data with distinct boundaries, enabling pre-training across domains for better generalization. Importantly, these unified graph features can represent data from any domain.
@@ II-D2 The Refinement of Cross-domain Difficulty
    High-Difficulty (When $|D_{T_{j}}\cap D_{S_{i}}|=0$): In this case, the source and target domains share no common data or features, indicating a complete lack of overlap. This extreme training setting is often adopted when the target domain has very limited data. It also aligns with the ultimate goal of CDGL—achieving open cross-domain graph learning under challenging conditions. A model that performs well here typically demonstrates strong generalizability.
@@
    Moderate-Difficulty (When $0<|D_{T_{j}}|-|D_{T_{j}}\cap D_{S_{i}}|<|D_{T_{j}}|$): In this scenario, the target domain shares either partial or complete overlap with the source domain. The difficulty is moderate because CDGL models can learn shared cross-domain knowledge in advance, reducing the overall challenge. However, unique domain characteristics may still require fine-tuning or domain-specific adjustments to address discrepancies.
@@
    Low-Difficulty (When $|D_{T_{j}}|-|D_{T_{j}}\cap D_{S_{i}}|=0$): This scenario indicates that the target domain is entirely equal to the source domain, with both domains sharing the same representation space. Cross-domain learning in this setting is relatively easy because the model does not need to transfer knowledge from source to target domains but only needs to capture cross-domain knowledge directly. This situation is more suitable for cases where both domains have abundant data, and the goal is to use a single model to handle tasks across different domains. Note that, in general, high overlap does not necessarily guarantee good generalization.
@@ III Structure-oriented Approach
Structural information, as the distinctive and core characteristic of graph data, is essential for addressing real-world graph-related problems. Existing graph models have demonstrated the ability to effectively capture structural information in single-domain scenarios. However, in cross-domain scenarios, the graphs from different domains exhibit varying node connectivity patterns, with some being densely connected and others more sparsely connected. These structural differences pose significant challenges to current graph models. To address this, structure-oriented approaches aim to improve the performance of existing models by extracting commonalities in structural information across domains. Based on how structural information is utilized, these approaches can be broadly categorized into three main branches: Structure Generation and Structure Contrast.
Figure 3: The core ideas behind structure-oriented approaches for capturing cross-domain knowledge consist of two main strategies: a) structure generation approaches, which employ a shared generator module to generate new and distinct structures for graph data from both source and target domains; b) structure contrast approaches, which construct positive and negative sample pairs for various graph data or leverage graph structure information to train GNNs and then make model parameter contrast across domains.
III-A Structure Generation
To effectively capture structural information, a useful approach is to establish a unified node connection strategy for graphs, ensuring that the graph data in both source domains and target domains follow the same connection rules. This allows the graph data from different domains to share common structural characteristics. The CDGL method based on structure generation is proposed with this idea in mind. It employs a shared module to generate new and distinct structures for graph data from both source and target domains. These newly generated structures enable the graph model to adopt a unified perspective for analyzing structural information across various domains, thereby enhancing the model’s performance on different graph-related tasks, as shown in Figure 3. For instance, GraphControl [22] introduces ControlNet, designed specifically for graph data, to extract downstream-specific graph structures from target domain graphs. Finally, by incorporating relevant feature information into the extracted subgraph data, cross-domain learning is achieved through fine-tuning or prompting. UniAug [25] is the first to introduce diffusion models for CDGL. Training a graph diffusion model on multi-source graph structure data enables the generation of transferable graph structures tailored to various downstream domains. GA2[23] is the first to reformulate graph data from different domains into subgraphs. In the pretraining phase, it leverages an autoencoder architecture to embed and reconstruct these subgraphs, enabling the graph model to utilize diverse structural information. OpenGraph [24] carefully a structure-based node tokenization scheme for graph domains, then introduces a scalable graph transformer to learn the node token embeddings and subsequently employs the LLM techniques to generate different graph structures, finally tackle various downstream graphs with variations in node sets and feature semantics.
III-B Structure Contrast
In contrast to structure generation, the goal of structure contrast is to effectively compare the structures of different graph datasets and identify common structural patterns across domains, as shown in Figure 3. This approach centers on designing negative and positive sample pairs to enable meaningful comparisons between graph structures, or on learning a set of weights to compare different structure-related model parameters. For instance, GCC [26] employs structural contrastive learning to pre-train graph neural networks in a cross-domain manner, enabling models to learn transferable structural representations across diverse domains. PCRec [27] pre-trains a graph neural network using self-supervised contrastive learning on source domain graphs to capture transferable structural information, then transfers the pre-trained encoder to initialize node embeddings on the target domain, enhancing recommendation performance by balancing source and target information while reducing bias from source domains. GRADE [28] introduces a graph adaptive network for CDGL by measuring and minimizing structural distribution shifts between source and target graphs using a Graph Subtree Discrepancy based on the Weisfeiler-Lehman graph isomorphism test, effectively comparing structural information for improved transferability. FedStar [30] leverages federated learning to train structure-only graph models on different clients, compares their parameters, and integrates them via weighted averaging to create a global model that merges cross-domain structural information. APT [29] leverages structural properties to contrast graph data during cross-domain pretraining, developing graph selection strategies that help the pretrained model more effectively capture cross-domain structural patterns. BooG [31] boosts the cross-domain learning of graph foundation models by unifying structural characteristics through virtual super nodes and edges, and employs a novel contrastive learning pre-training objective to learn expressive, generalizable graph representations. ProCom [32] enhances cross-domain targeted community detection by capturing structural information of latent communities through a dual-level context-aware pre-training method and a targeted community-guided prompting mechanism. RiemannGFM [33] captures cross-domain knowledge by learning a universal structural vocabulary of graph substructures—such as trees and cycles—embedding them into Riemannian manifolds and leveraging geometric contrastive learning for robust and transferable representations.
@@ III-C Discussions
Structure-oriented approaches have demonstrated superior performance on realizing CDGL, being able to handle various complex graph structural information. Moreover, they, specifically the structure generation approaches, also exhibit strong flexibility, as the unified node connection rules can generate meaningful structures for the graphs from unseen domains. Another advantage of these methods is their ability to explore the potential for uncovering commonalities across graph data from different domains, with structural information serving as the foundation of such data. For graph data, while structural information forms the foundation, it is the feature information that plays a crucial role in conveying substantive meaning. However, structure-oriented approaches mainly focus on capturing structural information but sometimes ignore the feature information. Taking structure generation approaches as an example, they need to consider the coherence between the generated structure and feature information to ensure effective communication across graph data from different domains.
@@ III-D Significance and Future Implications
Figure 4: The core ideas of feature-oriented approaches: a) feature align approaches, which directly leverage a trainable feature extractor to mine the commonalities among source and target domains; b) feature misalign approaches, which need to align the graph data with different feature dimensions across domains for training a general graph models. Note that both approaches can be further enhanced by prompting techniques.
@@ IV Feature-oriented Approach
As discussed in Section III-C, feature information plays a vital role in conveying substantive meaning in graph data, helping to differentiate between domains. The essence of feature-oriented approaches is to integrate feature information from various domains to discover commonalities across cross-domain features. However, training a general graph model on data from multiple domains presents unique challenges, primarily due to the feature differences across domains, particularly when feature dimensions do not align. For example, in most cases, if the feature dimensions of source domain $A$ and target domain $B$ differ, their semantic meanings will inevitably differ as well, since it is highly unlikely that domains with different dimensions share the same semantics. On the other hand, when A and B have the same feature dimensions, the semantics may either be identical or distinct. If the semantics are aligned, no additional adjustments are needed, and the domains can essentially be treated as one. However, when the semantics differ, further steps are required to bridge the semantic gap. Therefore, in this section, we broadly categorize models into Dimension Align and Dimension Misalign approaches, based on whether they require feature dimension alignment.
IV-A Dimension Align
The earliest feature-oriented approaches are Dimension Align, meaning they perform CDGL when the feature dimensions of the graph data in the source and target domains are consistent. Typically, when the feature dimensions align, the source and target domains share a high-level domain. Dimension Align approaches are generally considered part of limited CDGL. For instance, KTN [34] introduces different feature extractors for different node types in heterogeneous graph neural networks, and then uses domain adaptation strategies to integrate cross-domain feature information. NaP [35] improves out-of-distribution generalization in graph contrastive learning by dynamically treating semantically similar cross-domain negative pairs as positive pairs based on embedding similarity, reducing the distribution gap between domains, achieving CDGL to some extent. GPF-plus [36] and RELIEF [37] are two newly proposed feature-based graph prompt techniques to align the pre-training knowledge to different downstream datasets, achieving limited CDGL across molecular domains, where the dimension of different molecular datasets can be represented in the same feature space [88].
IV-B Dimension Misalign
Since Dimension Align approaches are significantly limited in cross-domain scale compared to open CDGL, which has already been widely achieved in NLP and CV thanks to the rapid development of LLMs, Dimension Misalign approaches have been proposed in recent years. These methods can map graph features from different domains into a shared embedding space with the same dimension. By aligning graph feature dimensions, they relax the limitation of cross-domain scale. Depending on the feature mapping strategy used, these approaches are classified as either conditional or open CDGL. For example, OFA [38] and ZeroG [41] both use LLMs to align text-attributed features from different domains. OFA freezes LLMs and directly trains a Multi-layer Perceptron (MLP) on aligned node features to achieve zero-shot node classification, while ZeroG [41] further utilizes a LoRA fine-tuning strategy to align semantic information across different domains and then leverages a text similarity task to conduct node classification results, achieving CDGL. Similar to the former approaches, GraphAlign [39] also uses LLMs to algin features, while it additionally designs alignment strategies of feature encoding, normalization, alongside a mixture-of-feature-expert module. Besides, CDFS-GAD [40] employs a domain-adaptive graph contrastive learning module that enhances cross-domain feature alignment by optimizing contrastive loss to bring similar nodes closer and push dissimilar ones apart. Additionally, a domain-specific prompt-tuning module customizes the encoder to effectively capture and retain both common and unique features across domains. Furthermore, DAGPrompT [42] proposes GLoRA to adapt feature projection through low-rank matrix updates, aligning pre-trained GNNs with downstream task distributions.

### v2

@@ II-D Cross-domain graph learning
Definitions. Derived from cross-domain learning, CDGL aims to capture cross-domain knowledge from given $m$ source graph domains $\mathbb{D}_{S}=\{\mathcal{D}_{G_{S_{1}}},\dots,\mathcal{D}_{G_{S_{i}}}|i=1,\dots,m\}$ and then transfer the captured knowledge to learn a set of task functions (e.g., node classification, graph classification, etc.) on $n$ target graph domains $\mathbb{D}_{T}=\{\mathcal{D}_{G_{T_{1}}},\dots,\mathcal{D}_{G_{T_{i}}}|i=1,\dots,n\}$.
@@ II-D1 The Refinement of Cross-domain Scale
    Limited cross-domain graph learning refers to scenarios where graph samples from source and target domains share a set of similar or consistent attributes (i.e., from the same field). For example, NaP [wang2024NAP] divides the Facebook100 dataset—which contains social networks from 100 U.S. universities—into different domains, treating the social network of each university as a unique domain. The data from each domain share similar attributes, except for the university types. CDGL is then conducted across different universities. Similarly, SDGA [qiao2023SDGA] treats ACMv9, Citationv1, and DBLPv7—datasets constructed from ArnetMiner [tang2008Arnetminer]—as distinct domains for cross-domain learning.
@@
    Conditional cross-domain graph learning refers to situations where the boundaries between source domains and graph domains are clearly defined, but their attributes can be transformed into a common format (e.g., text) and covered by a specific high-level field. For example, OFA [liu2024OFA] can convert cross-domain graph data into a text-attribute graph (TAG) format. However, it faces difficulties when applied to non-TAG domains, such as brain networks [yang2022METABrainC].
@@
    Open cross-domain graph learning refers to scenarios where graph data from both source and target domains can be arbitrary and do not need to satisfy specific conditions. For example, GCC [qiu2020gcc] focuses on leveraging only the structural information from datasets across domains, using it to pre-train models and learn common structural patterns in graph data. GCOPE [zhao2024GCOPE] presents a general framework for cross-domain graph pre-training that unifies the feature dimensions of graph data with distinct boundaries, enabling pre-training across domains for better generalization. Importantly, these unified graph features can represent data from any domain.
@@ II-D2 The Refinement of Cross-domain Difficulty
    High-Difficulty (When $|\mathbb{D}_{T}\cap\mathbb{D}_{S}|=0$): In this case, the source and target domains share no common data or features, indicating a complete lack of overlap. This extreme training setting is often adopted when the target domain has very limited data. It also aligns with the ultimate goal of CDGL—achieving open cross-domain graph learning under challenging conditions. A model that performs well here typically demonstrates strong generalizability.
@@
    Moderate-Difficulty (When $0<|\mathbb{D}_{T}|-|\mathbb{D}_{T}\cap\mathbb{D}_{S}|<|\mathbb{D}_{T}|$): In this scenario, the target domain shares either partial or complete overlap with the source domain. The difficulty is moderate because CDGL models can learn shared cross-domain knowledge in advance, reducing the overall challenge. However, unique domain characteristics may still require fine-tuning or domain-specific adjustments to address discrepancies.
@@
    Low-Difficulty (When $|\mathbb{D}_{T}|-|\mathbb{D}_{T}\cap\mathbb{D}_{S}|=0$): This scenario indicates that the target domain is entirely equal to the source domain, with both domains sharing the same representation space. Cross-domain learning in this setting is relatively easy because the model does not need to transfer knowledge from source to target domains but only needs to capture cross-domain knowledge directly. This situation is more suitable for cases where both domains have abundant data, and the goal is to use a single model to handle tasks across different domains. Note that, in general, high overlap does not necessarily guarantee good generalization.
@@ III Structure-oriented Approach
Structural information, as the distinctive and core characteristic of graph data, is essential for addressing real-world graph-related problems. Existing graph models have demonstrated the ability to effectively capture structural information in single-domain scenarios. However, in cross-domain scenarios, the graphs from different domains exhibit varying node connectivity patterns, with some being densely connected and others more sparsely connected. These structural differences pose significant challenges to current graph models. To address this, structure-oriented approaches aim to improve the performance of existing models by extracting commonalities in structural information across domains. Based on how structural information is utilized, these approaches can be broadly categorized into two main branches: Generative and Contrastive.
III-A Generative
To effectively capture structural information, a useful approach is to establish a unified node connection strategy for graphs, ensuring that the graph data in both source domains and target domains follows the same connection rules. This allows the cross-domain graph data to share the same structural space. Structure-oriented generative CDGL methods are developed based on this principle. It employs a shared module to generate new and distinct structures for graph data from both source and target domains. These newly generated structures enable the graph model to adopt a unified perspective for analyzing structural information across various domains, thereby enhancing the model’s performance on different graph-related tasks. As shown in Figure 3, generative methods typically begin by employing a structure generation module $Generator(\cdot)$ to edit and transform the input graph $\mathcal{G_{S}}\in\mathbb{D}_{G_{S}}$, resulting in new graph structures and thereby forming a modified source domain $\mathbb{D}^{{}^{\prime}}_{G_{S}}$. A graph model $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is then trained on $\mathbb{D}^{{}^{\prime}}_{G_{S}}$ and subsequently generalized to the target domain $\mathbb{D}_{T}$ via a task-specific prediction head $TaskH(\cdot)$, which outputs the final prediction results:
Structure Generation: $\mathcal{G}^{{}^{\prime}}_{S}=Generator(\mathcal{V}_{S},\mathcal{E}_{S},\mathcal{A}_{S}),$ (5)
Generalization: $\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{{}^{\prime}}_{T})),$ (6)
Here, $Generator(\cdot)$ aims to generate unified structural patterns across domains, and $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is expected to generalize to $\mathbb{D}_{T}$ despite domain-specific differences.
For instance, GraphControl [zhu2024graphcontrol] utilizes structure pretraining on source graphs to learn common structural patterns. Inspired by ControlNet, it generates a new structure from target node features and combines it with the original graph structure. This feature-derived structure is injected into a trainable copy of the pretrained model through zero-MLPs, allowing the model to adapt to the target domain and effectively transfer structural knowledge. UniAug [tang2024UniAug] introduces diffusion models into cross-domain graph learning for the first time. It pre-trains a graph diffusion model on source domain structures to learn shared structural patterns. Then, it modifies the target domain graphs using the frozen pre-trained diffusion model to generate new structures, which are used for fine-tuning or prompting a downstream classifier, thereby achieving structure-level transfer across domains. GA2[hu2024GA2E] is the first to reformulate graphs from different domains into subgraphs for cross-domain learning. It leverages an autoencoder framework during pretraining to embed and reconstruct these subgraphs, allowing the model to capture and unify diverse structural patterns. This subgraph-based reconstruction enables effective structural knowledge transfer to target domains. OpenGraph [xia2024opengraph] first designs a node tokenization scheme that converts nodes from various source domain graphs into a unified token set based on their structural information (i.e., adjacency matrices). A scalable Transformer is then pre-trained on these tokens to learn a shared representation space. To mitigate limited structural coverage in the training data, OpenGraph uses LLMs to synthesize entirely new node–edge sets, broadening the structural distribution. The pre-trained Transformer captures shared structural patterns that can be effectively transferred to target domains. GFT [wang2024gft] transforms graphs into computation trees derived from GNN message passing, then generates a discrete tree vocabulary through self-supervised learning, embedding structural commonalities across domains. During fine-tuning, downstream tasks are reformulated as tree classification using this shared vocabulary, enabling efficient cross-domain transfer. GFSE [chen2025gfse] introduces four structure-aware self-supervised tasks (e.g., Shortest Path Distance Regression [Dijkstra1959shortestpath1, li2020shortestpath2]) during pre-training to uncover structural commonalities across multi-domain graph datasets. In particular, the Shortest Path Distance Regression task dynamically extracts and models diverse shortest-path substructures during the pre-training phase, thereby enabling the pre-trained model to understand cross-domain graph structures.
Figure 3: Structure-oriented CDGL: (a) a shared generator produces diverse structures for both source and target domains; (b) structure contrast explicitly or implicitly constructs positive–negative pairs in the source domain to improve generalization to target domains.
III-B Contrastive
In contrast to structure generation, structure-oriented contrastive methods aim to compare structural patterns across different graph datasets and identify domain-invariant commonalities among them. As shown in Figure 3, contrastive methods typically begin by employing a structure contrastive module $Contrast(\cdot)$ to compare and modify the input graph $\mathcal{G_{S}}\in\mathbb{D}_{G_{S}}$, which can produce positive and negative sample pairs $(\mathcal{G}^{pos}_{S},\mathcal{G}^{neg}_{S})$. A graph model $\mathcal{M}_{G_{S}}$ is then trained on the produced sample pairs and subsequently generalized to the target domain $\mathbb{D}_{T}$ via a task-specific prediction head $TaskH(\cdot)$, which outputs the final prediction results:
Structure Contrast: $(\mathcal{G}^{pos}_{S},\mathcal{G}^{neg}_{S})=Contrast(\mathcal{V}_{S},\mathcal{E}_{S},\mathcal{A}_{S}),$ (7)
Generalization: $\hat{y}_{T}=TaskH(\mathcal{M}_{G_{S}}(\mathcal{G}_{T})),,$ (8)
Here, $Contrast(\cdot)$ is designed to perform structural comparative analysis across domains, facilitating the discovery of domain-invariant patterns. $\mathcal{M}_{G_{S}}$ is expected to generalize to $\mathbb{D}_{T}$ despite domain-specific differences.
For instance, GCC [qiu2020gcc] leverages two structural augmentations to generate numerous subgraph instances from source domains. Positive pairs are formed from instances of the same graph, while negative pairs come from different graphs. These pairs are used to pretrain GNNs with a contrastive objective, enabling the model to capture structural commonalities in source domains and transfer them to target domains via fine-tuning or prompting. PCRec [wang2021pcrec] constructs positive and negative sample pairs by generating two augmented views of user-item graphs (without node features) in source domains, and uses contrastive learning to capture common structural patterns. It then transfers the pre-trained graph encoder to the target domain by fine-tuning it on target data, improving recommendation performance. GRADE [wu2023GRADE] introduces a Graph Subtree Discrepancy (GSD) metric, inspired by the Weisfeiler-Lehman test, to measure differences in structure between source and target graphs. Then, it trains a graph model by minimizing both the widely used task-specific loss on the source graph and this GSD term, which implicitly treats similar subtrees as positive pairs (pulled together) and dissimilar ones as negative pairs (pushed apart), enabling the trained model to capture the cross-domain structural commondalities across domains and be generalized to target domains effectively. FedStar [tan2023FedStar] leverages federated learning to capture shared structural patterns across domains. Specifically, each client locally trains a structural encoder to capture its domain-specific graph structures. Then, the server aggregates these encoders’ parameters into a unified global model through weighted averaging. This iterative process implicitly reinforces structural patterns common to multiple domains (pulling ”positive samples” closer) while suppressing domain-specific structural features (pushing ”negative samples” apart), thereby achieving an implicit cross-domain structural contrastive transfer. APT [xu2023APT] proposes a graph selection strategy based on structural properties such as average degree and network entropy, aiming to select the most informative source graphs for contrastive pre-training. By constructing positive and negative sample pairs from these representative graphs, the model can be pre-trained more effectively to capture structural commonalities across domains and improve generalization to target graphs. BooG [cheng2024BooG] boosts the cross-domain learning of graph foundation models by unifying structural characteristics through virtual super nodes and edges, and employs a novel contrastive learning method to construct positive and negative sample pairs and pre-train a graph model, thereby capturing cross-domain structure commonalities that can be generalized to target domains. ProCom [wu2024procom] introduces a dual‐level contrastive pre‐training (i.e., node‐to‐context and context‐to‐context levels) method to learn shared community structures, thereby improving the generalizability of pre-trained GNN-based community detection models to various target domains. RiemannGFM [sun2025RiemannGFM] captures cross-domain knowledge by learning a universal structural vocabulary of graph substructures (e.g., trees and cycles), embedding them into Riemannian manifolds, and finally leveraging geometric contrastive learning for pre-training a graph model that can be easily transferred to target domains.
@@ III-C Discussions
Structure-oriented approaches have demonstrated superior performance on realizing CDGL, being able to handle various complex graph structural information. Moreover, they, specifically the generative approaches, also exhibit strong flexibility, as the unified node connection rules can generate meaningful structures for the graphs from unseen domains. Another advantage of these methods is their ability to explore the potential for uncovering commonalities across graph data from different domains, with structural information serving as the foundation of such data. For graph data, while structural information forms the foundation, it is the feature information that plays a crucial role in conveying substantive meaning. However, structure-oriented approaches mainly focus on capturing structural information but sometimes ignore the feature information. Taking structure generation approaches as an example, they need to consider the coherence between the generated structure and feature information to ensure effective communication across graph data from different domains.
@@ IV Feature-oriented Approach
As discussed in Section III-C, feature information plays a vital role in conveying substantive meaning in graph data, helping to differentiate between domains. The essence of feature-oriented approaches lies in learning transferable feature representations by identifying feature-level commonalities across domains. However, training a general graph model on data from multiple domains poses significant challenges, primarily due to differences in feature semantics and dimensions. To address these challenges, existing feature-oriented CDGL approaches can be categorized into two groups: In-space and Cross-space, depending on whether they require alignment of feature dimensions. In most cases, if the feature dimensions of source domain $A$ and target domain $B$ differ, their semantic meanings will inevitably differ as well, since it is highly unlikely that domains with different dimensions share the same semantics—this scenario typically requires Cross-space approaches. On the other hand, when $A$ and $B$ have the same feature dimensions, the semantics may either be identical or distinct. If the semantics are aligned, no additional adjustments are needed, and the domains can essentially be treated as one. However, when the semantics differ, further steps are required to bridge the semantic gap—this case calls for In-space approaches.
IV-A In-space
The earliest feature-oriented approaches are In-space, meaning they perform CDGL when the feature dimensions of the graph data in the source and target domains are consistent, primarily focusing on aligning the feature semantic information. As shown in Figure 4, in-space methods typically begin by employing a semantic alignment module $SAlign(\cdot)$ to map the input graph feature $\mathcal{X}_{S}\in\mathbb{D}_{G_{S}}$ into a new feature $\mathcal{X}^{{}^{\prime}}_{S}$, thereby resulting in new source domains $\mathbb{D}^{{}^{\prime}}_{G_{S}}$. A graph model $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is then trained on the new source domains and subsequently generalized to the target domains $\mathbb{D}_{G_{T}}$ via a task-specific prediction head $TaskH(\cdot)$, which outputs the final prediction results. Besides, a promoting module $prompt(\cdot)$ is optionally leveraged to further enhance the cross-domain transferability:
Semantic Alignment: $\mathcal{X}^{{}^{\prime}}_{S}=SAlign(\mathcal{X}_{S}),$ (9)
Generalization: $\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{{}^{\prime}}_{T})),$ (10)
(Prompt): $\mathcal{X}^{p}_{T}=prompt(\mathcal{X}_{T}),$ (11)
$\mathcal{X}^{p^{{}^{\prime}}}_{T}=SAlign(\mathcal{X}^{p}_{T}),$ (12)
$\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{p^{{}^{\prime}}}_{T}))$ (13)
Here, $SAlign(\cdot)$ is designed to align feature semantics across domains. $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is expected to generalize to $\mathbb{D}_{T}$ despite semantic-specific differences. $prompt(\cdot)$ is used to modify $\mathcal{X}_{T}$ into $\mathcal{X}^{p}_{T}$ to better align with the semantic space captured by $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$, thereby facilitating cross-domain transfer.
Note that the source and target domains typically share a high-level domain (e.g., molecular domain [li2022molecularsmall]) when the feature dimensions align. Thus, in-space approaches are generally considered part of the limited CDGL. For instance, KTN [yoon2022KTN] introduces a trainable knowledge transfer network that models the theoretical transformation between feature extractors of different node types in a heterogeneous graph. Specifically, it learns to map the target node embeddings into the source domain feature space using adjacency matrices and a shared learnable transformation matrix, thereby aligning their semantic representations and enabling effective cross-domain feature semantic transfer. NaP [wang2024NAP] proposes a contrastive learning framework that dynamically re-labels semantically similar cross-domain negative pairs as positive ones based on embedding similarity. By doing so, it pre-trains a domain-invariant GNN encoder that implicitly captures shared feature-level semantic commonalities across source domains and reduces the distribution discrepancy. The pre-trained encoder can be directly generalized to target domains, achieving cross-domain feature semantic transfer. GPF-plus [fang2024gpfplus] and RELIEF [zhu2024relief] are graph prompt-based methods that introduce learnable prompt features to augment graph representations. By adding prompt tokens to the original node features in target domains, they guide the GNN encoder to align input representations with a pretrained semantic space. These prompts implicitly capture semantic commonalities learned from source domains and enable effective cross-domain semantic transfer. DAGPrompT [chen2025DAGPrompT] proposes a GLoRA module, inspired by the success of LoRA in NLP [hu2022loraInNLP], to achieve cross-domain semantic alignment through low-rank matrix updates for adapting the feature semantics of target domains to the pre-trained GNNs on source domains. GADT3 [pirhayatifard2025crossdomainGADT3] leverages test-time training to achieve cross-domain adaptation, where a homophily-based self-supervised loss guides a target-specific encoder to align domain-invariant node patterns with a frozen source decoder, thereby transferring anomaly-discriminative knowledge without access to source data or target labels.
Figure 4: Feature-oriented CDGL: a) tackles the feature semantic alignment across domains; b) further aligns the feature dimensions across various graph domains to achieve more general CDGL. Prompt techniques are optional for both approaches.
IV-B Cross-space
Since in-space approaches are significantly limited in cross-domain scale compared to open CDGL, Cross-space approaches have been proposed in recent years to tackle the feature dimension misalignment problem. Specifically, these methods can map cross-domain graph features into a shared embedding space with the same dimension, which enables the graph model to take various cross-domain graphs as input, thereby learning shared feature patterns across domains and relaxing the limitation of cross-domain scale. As shown in Figure 4, cross-space methods typically begin by employing a dimension alignment module $DAlign(\cdot)$ to map the input graph features from different domains with different dimensions, denoted as $\mathcal{X}_{S_{1}},\mathcal{X}_{S_{2}}\in\mathbb{D}_{G_{S}}$, $\mathcal{X}_{S_{1}}\in\mathbb{R}^{D}$ and $\mathcal{X}_{S_{2}}\in\mathbb{R}^{\hat{D}}$, into new ones $\mathcal{X}^{{}^{\prime}}_{S_{1}},\mathcal{X}^{{}^{\prime}}_{S_{2}}\in\mathbb{R}^{S}$, thereby resulting in new source domains $\mathbb{D}^{{}^{\prime}}_{G_{S}}$. A graph model $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is then trained on the new source domains and subsequently generalized to the target domain $\mathbb{D}_{T}$ via a task-specific prediction head $TaskH(\cdot)$, which outputs the final prediction results. Besides, a promoting module $prompt(\cdot)$ is also optional:
Dimension Alignment: $\mathcal{X}^{{}^{\prime}}_{S_{1}}=DAlign(\mathcal{X}_{S_{1}}),$ (14)
$\mathcal{X}^{{}^{\prime}}_{S_{2}}=DAlign(\mathcal{X}_{S_{2}}),$ (15)
Generalization: $\mathcal{X}^{{}^{\prime}}_{T}=DAlign(\mathcal{X}_{T}),$ (16)
$\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{{}^{\prime}}_{T})),$ (17)
(Prompt): $\mathcal{X}^{p}_{T}=prompt(\mathcal{X}^{{}^{\prime}}_{T}),$ (18)
$\hat{y}_{T}=TaskH(\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}(\mathcal{G}^{p}_{T}))$ (19)
Here, $DAlign(\cdot)$ is designed to align feature dimensions across domains. $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$ is expected to generalize to $\mathbb{D}_{T}$ despite dimension-specific differences. $prompt(\cdot)$ is used to modify $\mathcal{X}_{T}$ into $\mathcal{X}^{p}_{T}$ to better align with the feature space captured by $\mathcal{M}_{\mathbb{D}^{{}^{\prime}}_{G_{S}}}$.
For instance, OFA [liu2024OFA] and ZeroG [li2024zerog] both leverage LLMs to align text-attributed feature dimensions across domains. Particularly, OFA [liu2024OFA] freezes an LLM to generate LLM-based embeddings as node features for both source and target graphs, which are then fed into a lightweight classification head trained on source domains, enabling few-shot or zero-shot node classification on unseen target domains. Compared with OFA, ZeroG [li2024zerog] executes parameter-efficient fine-tuning (LoRA) on a frozen LLM, aligning node attribute and class description embeddings from source domains into a shared feature space. Then, it further enriches LLM-based node features by a semantic prompt node, and finally performs zero-shot node classification by measuring text similarity between node embeddings and class embeddings. Similar to the former approaches, GraphAlign [hou2024graphalign] also applies LLMs to align node feature dimensions, while additionally designing alignment strategies for feature encoding and normalization, alongside a mixture-of-feature-expert module for better handling cross-domain semantic information beyond dimension differences across domains to transfer feature commonalities. CDFS-GAD [chen2024CDFSGAD] employs a domain-adaptive graph contrastive learning module to pre-train a graph model, which can align the dimensions of graphs from various source domains and then construct positive and negative sample pairs across domains. Then, a domain-specific prompt-tuning module is introduced to align the features of target domains with pre-trained feature representation space while preserving their unique features. UNPrompt [niu2024UNPrompt] first applies feature projection (e.g., SVD [stewart1993SVD]) to align node feature dimensions across graph domains, then injects a small set of learnable prompt tokens into node attributes. These tokens are trained on a single source dataset and then frozen, enabling zero-shot anomaly detection by aligning the semantic space of target graphs with the pre-trained source representation.